# On the existential theory of the completions of a global field

### v2

@@ 6. Existential $\mathfrak{L}_{\rm val}$-theory with parameters
    Every field $L$ which is existentially closed in $L(\!(t)\!)$ is existentially closed in every extension $F/L$ for which there exists a valuation $v$ on $F$ trivial on $L$ with residue field $Fv=L$.
@@ 1. UR($t$)
If $t$ does not have a $p$-th root in $Fv$, and hence no $p$-th root in $F^{\prime}v^{\prime}$, then $Fv/C$ and $F^{\prime}v^{\prime}/C$ are separable, and in this case the claim follows from [ADF23, Corollary 4.16] (using (R4)). So let us suppose that $t$ is a $p$-th power in $Fv$ and therefore also in $F^{\prime}v^{\prime}$. The identity on the perfect field $C_{0}$ extends to sections $\zeta\colon Fv\to F$ and $\zeta^{\prime}\colon F^{\prime}v^{\prime}\to F^{\prime}$ of the residue homomorphisms [ADF23, Proposition 4.5]. Let $s:=\zeta(t^{1/p})$ and $s^{\prime}:=\zeta^{\prime}(t^{1/p})$, so that $v(t-s^{p})>0$, $v^{\prime}(t-s^{\prime p})>0$. By UR($t$), $\pi:=t-s^{p}$ and $\pi^{\prime}:=t-s^{\prime p}$ are uniformisers for $v$ and $v^{\prime}$, respectively. As $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, the relative algebraic closures $D$ and $D^{\prime}$ of $C$ in $Fv$ respectively $F^{\prime}v^{\prime}$ are isomorphic (see again [FJ08, Lemma 20.6.3(b)]), and we fix once and for all an isomorphism and identify them. As $D$ has imperfect exponent at most $1$ (since it can be no higher than that of $C$) the extensions $Fv/D$ and $F^{\prime}v^{\prime}/D$ are separable [FJ08, Lemma 2.7.5]. The $C_{0}$-isomorphism $\zeta^{\prime}\circ\zeta^{-1}|_{\zeta(D)}\colon\zeta(D)\rightarrow\zeta^{\prime}(D)$ extends uniquely to an isomorphism $g\colon\zeta(D)(\pi)\rightarrow\zeta^{\prime}(D)(\pi^{\prime})$ with $g(\pi)=\pi^{\prime}$. As $v$ is trivial on $\zeta(D)$ and $v(\pi)>0$, $v|_{\zeta(D)(\pi)}$ is the $\pi$-adic valuation on the function field $\zeta(D)(\pi)$, and similarly for $v^{\prime}|_{\zeta^{\prime}(D)(\pi^{\prime})}$. Therefore, $g$ is in fact an isomorphism of valued subfields of $(F,v)$ respectively $(F^{\prime},v^{\prime})$. As $g(s)=s^{\prime}$ and $g(\pi)=\pi^{\prime}$, we have $g(t)=t$, so $g$ is even a $C_{0}(t)$-isomorphism.
Now [ADF23, Proposition 4.11] implies (using (R4)) that the universal/existential $\mathfrak{L}_{\rm val}$-theories of $(F,v)$ and $(F^{\prime},v^{\prime})$ agree, with parameters from $\zeta(D)(\pi)$ and $\zeta^{\prime}(D)(\pi^{\prime})$ respectively, where the parameter fields are identified via $g$. Let $c\in C$ be arbitrary, and let $f\in C_{0}(t)[x]$ be its minimal polynomial over $C_{0}$. Now $c$ is the unique zero of $f$ in $F$ with the same residue as $\zeta(c)\in\zeta(D)$. This identifies $c$ as the unique witness in $(F,v)$ of the quantifier-free $\mathfrak{L}_{\rm val}(\zeta(D)(\pi))$-formula
@@
$0\to\mathfrak{m}/\mathfrak{m}^{2}\to Fv\otimes_{\mathcal{O}}\Omega_{\mathcal{O}}\to\Omega_{Fv}\to 0$
@@
This could be deduced from the characterisation given in Remark 6.3, but we give a direct argument. Assume that $v$ does not ramify over $k(t)$. If $v(t)<0$, this implies that $v(t)=-1$, and so for arbitrary $s\in\hat{K}_{v}$ we have $v(t-s^{p})=\min(v(t),v(s^{p}))\leq-1$ since $v(t)\neq v(s^{p})$. Assume now that $v(t)\geq 0$. The restriction $v|_{k(t)}$ is the place of $k(t)$ associated to an irreducible polynomial $f\in k[t]$. Since $k[t]/(f)$ is perfect, there exists $g\in k[t]$ such that $f|t-g^{p}$. As the right-hand side is a separable polynomial, $f^{2}\nmid t-g^{p}$. Since $v$ does not ramify over $k(t)$, we deduce that $v(t-g^{p})=1$. For arbitrary $s\in\hat{K}_{v}$, we write $t-s^{p}=(t-g^{p})+(g-s)^{p}$, and observe that the two summands have different value under $v$, so $v(t-s^{p})=\min(v(t-g^{p}),v((g-s)^{p}))\leq v(t-g^{p})=1$. This shows that UR($t$) holds in $\hat{K}_{v}$. ∎
@@ 2. $(b)$
    The universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of all $\hat{K}_{v}$, $v\in\mathbb{P}_{K}$, is the universal/existential theory of henselian nontrivially valued fields $(L,w)$ containing $K$ with residue field finite or pseudofinite such that UR($t$) holds unless $w(z_{u_{i}})>0$ for some $i=1,\dots,r$, and such that ${\rm Th}_{\forall/\exists}(L,w)={\rm Th}_{\forall/\exists}(\hat{K}_{v},\hat{v})$ whenever $w(z_{v})>0$ for some $v\in\mathbb{P}_{K}$, where for each $v\in\mathbb{P}_{K}$, $z_{v}\in K$ is chosen to have a zero only at $v$, and $u_{1},\dots,u_{r}$ is the list of places of $K$ which ramify over $k(t)$.
@@
Taking (a) into account, part (b) asserts that a field $(L,w)$ satisfies the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of all $\hat{K}_{v}$, $v\in\mathbb{P}_{K}$, if and only if $(L,w)$ either satisfies the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of almost all $\hat{K}_{v}$, or ${\rm Th}_{\forall/\exists}(L,w)={\rm Th}_{\forall/\exists}(\hat{K}_{v},\hat{v})$ for some $v\in\mathbb{P}_{K}$. This is true on general grounds. Indeed, suppose that $(L,w)$ does not satisfy the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of almost all $\hat{K}_{v}$, so that $(L,w)\models\neg\varphi$ for some universal/existential $\mathfrak{L}_{\rm val}(K)$-sentence $\varphi$ which is true in all $\hat{K}_{v}$ except for finitely many exceptional places $v_{1},\dotsc,v_{n}$. If $(L,w)$ also does not satisfy the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of any individual $(\hat{K}_{v},\hat{v})$, we in particular have universal/existential $\mathfrak{L}_{\rm val}(K)$-sentences $\varphi_{1},\dotsc,\varphi_{n}$ such that $(L,w)\models\neg\varphi_{i}$ but $(\hat{K}_{v_{i}},\hat{v}_{i})\models\varphi_{i}$ for every $i=1,\dotsc,n$. Therefore $(L,w)\not\models\varphi\vee\varphi_{1}\vee\dotsb\vee\varphi_{n}$ even though $(\hat{K}_{v},\hat{v})$ satisfies the sentence on the right-hand side for every $v\in\mathbb{P}_{K}$, and so $(L,w)$ is not a model of the universal/existential theory of all $(\hat{K}_{v},\hat{v})$. Conversely, if $(L,w)$ satisfies either the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of almost all $\hat{K}_{v}$, or of one individual $\hat{K}_{v}$, it is clear that $(L,w)$ satisfies the common universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of all $\hat{K}_{v}$. ∎
@@
Fix a separating element $t\in K$ and let $P_{0}\subseteq\mathbb{P}_{K}$ be the finite set of places that ramify over $k(t)$. Similar to before, the decidability of the universal/existential $\mathfrak{L}_{\rm val}(K)$-theory of almost all $\hat{K}_{v}$ follows immediately from Proposition 6.1, Lemma 4.4 and Lemma 2.1(b). In order to decide whether a given universal/existential $\mathfrak{L}_{\rm val}(K)$-sentence $\varphi$, holds in all $\hat{K}_{v}$ first check whether it holds in almost all of them. If yes, then by Theorem 6.5 and the completeness theorem, there is a proof of $\varphi$ from the set of axioms for henselian nontrivially valued fields extending the trivially valued field $C=K$ satisfying UR($t$) and with residue field pseudofinite. As this set of axioms is computable (see in particular Lemma 4.4), the set of consequences is computably enumerable, and so by enumerating all proofs from these axioms, we will eventually find a proof of $\varphi$. Apart from possibly using UR($t$), which holds in $\hat{K}_{v}$ for all $v\notin P_{0}$ (Lemma 6.4), this proof will involve only finitely many constants from $K$ and only finitely many axioms of pseudofinite fields, and so we can compute a finite set $P_{1}\subseteq\mathbb{P}_{K}$ such that $\hat{K}_{v}\models\varphi$ for every $v\notin P_{0}\cup P_{1}$. To determine whether $\hat{K}_{v}\models\varphi$ for each of the finitely many $v\in P_{0}\cup P_{1}$, we use [ADF23, Theorem 4.12], which (again assuming (R4)) gives a uniform decidability of each individual ${\rm Th}_{\forall/\exists}(\hat{K}_{v},\hat{v})$. ∎
@@ Acknowledgements
A. F. would like to thank Sylvy Anscombe for helpful discussions at an early stage of this project. A. F. was funded by the Deutsche Forschungsgemeinschaft (DFG) - 404427454.

### v3

@@ 6. Existential $\mathfrak{L}_{\rm val}$-theory with parameters
    Every field $L$ which is existentially closed in L((t))L(\!(t)\!) is existentially closed in every extension $F/L$ for which there exists a valuation $v$ on $F$ trivial on $L$ with residue field $Fv=L$.
In this work we will use (R4) only through results in [ADF23] that are proven conditionally on it, and therefore we refer the reader to [ADF23, Proposition 2.3] for how (R4) would follow from resolution of singularities, and to [ADF23, Remark 2.4] for fields $L$ for which it is known to hold. We will make use in particular of the following two main results that are proven conditionally on (R4):
Theorem 6.1 ([ADF23, Proposition 4.11, Corollary 4.16]).
Assume (R4).
  1. $(a)$
    Let $(C,u)$ be an equicharacteristic valued field with $uC=\mathbb{Z}$ and $\mathcal{O}_{u}$ excellent, and let $\pi\in C$ with $u(\pi)=1$. Let $(F,v)$ and $(F^{\prime},v^{\prime})$ be henselian extension of $(C,u)$ such that $v(\pi)$ is minimal positive in $vF$ and $v^{\prime}(\pi)$ is minimal positive in $v^{\prime}F^{\prime}$, and such that $Fv/Cu$ and $F^{\prime}v^{\prime}/Cu$ are separable. If $Fv$ and $F^{\prime}v^{\prime}$ have the same existential $\mathfrak{L}_{\rm ring}(Cu)$-theory, then $(F,v)$ and $(F^{\prime},v^{\prime})$ have the same existential $\mathfrak{L}_{\rm val}(C)$-theory.
  2. $(b)$
    Let $C$ be a field and let $T$ be the theory of equicharacteristic nontrivially valued henselian fields $(F,v)$ extending the trivially valued field $C$, such that $Fv/C$ is separable. Then $T({\rm Th}_{\forall/\exists}(k))$ is $\exists$-$C$-complete (cf. Theorem 4.1) for every separable extension $k/C$.
@@ 1. UR($t$)
If $t$ does not have a $p$-th root in $Fv$, and hence no $p$-th root in $F^{\prime}v^{\prime}$, then $Fv/C$ and $F^{\prime}v^{\prime}/C$ are separable, and in this case the claim follows from Theorem 6.1(b) (using (R4)). So let us suppose that $t$ is a $p$-th power in $Fv$ and therefore also in $F^{\prime}v^{\prime}$. The identity on the perfect field $C_{0}$ extends to sections $\zeta\colon Fv\to F$ and $\zeta^{\prime}\colon F^{\prime}v^{\prime}\to F^{\prime}$ of the residue homomorphisms [ADF23, Proposition 4.5]. Let $s:=\zeta(t^{1/p})$ and $s^{\prime}:=\zeta^{\prime}(t^{1/p})$, so that $v(t-s^{p})>0$, $v^{\prime}(t-s^{\prime p})>0$. By UR($t$), $\pi:=t-s^{p}$ and $\pi^{\prime}:=t-s^{\prime p}$ are uniformisers for $v$ and $v^{\prime}$, respectively. As $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, the relative algebraic closures $D$ and $D^{\prime}$ of $C$ in $Fv$ respectively $F^{\prime}v^{\prime}$ are isomorphic (see again [FJ08, Lemma 20.6.3(b)]), and we fix once and for all an isomorphism and identify them. As $D$ has imperfect exponent at most $1$ (since it can be no higher than that of $C$) the extensions $Fv/D$ and $F^{\prime}v^{\prime}/D$ are separable [FJ08, Lemma 2.7.5]. The $C_{0}$-isomorphism $\zeta^{\prime}\circ\zeta^{-1}|_{\zeta(D)}\colon\zeta(D)\rightarrow\zeta^{\prime}(D)$ extends uniquely to an isomorphism $g\colon\zeta(D)(\pi)\rightarrow\zeta^{\prime}(D)(\pi^{\prime})$ with $g(\pi)=\pi^{\prime}$. As $v$ is trivial on $\zeta(D)$ and $v(\pi)>0$, $v|_{\zeta(D)(\pi)}$ is the $\pi$-adic valuation on the function field $\zeta(D)(\pi)$, and similarly for $v^{\prime}|_{\zeta^{\prime}(D)(\pi^{\prime})}$. Therefore, $g$ is in fact an isomorphism of valued subfields of $(F,v)$ respectively $(F^{\prime},v^{\prime})$. As $g(s)=s^{\prime}$ and $g(\pi)=\pi^{\prime}$, we have $g(t)=t$, so $g$ is even a $C_{0}(t)$-isomorphism.
Now Theorem 6.1(a) implies (using (R4)) that the universal/existential $\mathfrak{L}_{\rm val}$-theories of $(F,v)$ and $(F^{\prime},v^{\prime})$ agree, with parameters from $\zeta(D)(\pi)$ and $\zeta^{\prime}(D)(\pi^{\prime})$ respectively, where the parameter fields are identified via $g$. Let $c\in C$ be arbitrary, and let $f\in C_{0}(t)[x]$ be its minimal polynomial over $C_{0}$. Now $c$ is the unique zero of $f$ in $F$ with the same residue as $\zeta(c)\in\zeta(D)$. This identifies $c$ as the unique witness in $(F,v)$ of the quantifier-free $\mathfrak{L}_{\rm val}(\zeta(D)(\pi))$-formula
@@
$0\to\mathfrak{m}/\mathfrak{m}^{2}\to Fv\otimes_{\mathcal{O}}\Omega_{\mathcal{O}}\to\Omega_{Fv}\to 0.$
@@
This could be deduced from the characterisation given in Remark 6.4, but we give a direct argument. Assume that $v$ does not ramify over $k(t)$. If $v(t)<0$, this implies that v(t)=1v(t)=-1, and so for arbitrary $s\in\hat{K}_{v}$ we have $v(t-s^{p})=\min(v(t),v(s^{p}))\leq-1$ since $v(t)\neq v(s^{p})$. Assume now that $v(t)\geq 0$. The restriction $v|_{k(t)}$ is the place of $k(t)$ associated to an irreducible polynomial $f\in k[t]$. Since $k[t]/(f)$ is perfect, there exists $g\in k[t]$ such that $f|t-g^{p}$. As the right-hand side is a separable polynomial, $f^{2}\nmid t-g^{p}$. Since $v$ does not ramify over $k(t)$, we deduce that $v(t-g^{p})=1$. For arbitrary $s\in\hat{K}_{v}$, we write $t-s^{p}=(t-g^{p})+(g-s)^{p}$, and observe that the two summands have different value under $v$, so $v(t-s^{p})=\min(v(t-g^{p}),v((g-s)^{p}))\leq v(t-g^{p})=1$. This shows that UR($t$) holds in $\hat{K}_{v}$. ∎
@@ 2. $(b)$
[…]
@@
[…]
@@
[…]
@@ Acknowledgements
The authors would like to thank the referee for suggestions to improve the exposition of this work. A. F. would like to thank Sylvy Anscombe for helpful discussions at an early stage of this project. A. F. was funded by the Deutsche Forschungsgemeinschaft (DFG) - 404427454. Both authors would like to thank the Hausdorff Research Institute for Mathematics Bonn, funded by the DFG under Germany’s Excellence Strategy (EXC-2047/1 – 390685813), for its hospitality during the trimester programme “Definability, decidability, and computability” while this paper was revised.